OpenStudy (kamille):

I need to solve: sin(4x)=3cos(2x)

OpenStudy (kamille):

@hartnn maybe you could help me now?

hartnn (hartnn):

write sin 4x = 2 sin 2x cos 2x

OpenStudy (anonymous):

we know that\[\sin 2 \theta = 2\sin \theta \cos \theta\]

OpenStudy (kamille):

yeah, I did it

OpenStudy (kamille):

2 sin 2x cos 2x do i need to rewrite it again, using sin(2x) and cos(2x) again?

OpenStudy (anonymous):

so \[2\sin 2x\cos 2x=3\cos 2x\] \[\cos2x(2\sin 2x-3)=0\] \[\cos2x=0\] or \[2\sin 2x-3=0\]

OpenStudy (kamille):

didnt get :/ can you explain?

hartnn (hartnn):

which step you didn't get ? he factored out cos 2x after subtracting 3cos2x from both sides..

OpenStudy (kamille):

oh, understand now:) Thank you a lot:)

OpenStudy (kamille):

didnt understand how -3 appeared in the breckets, but I do now:) THANK YOU A LOT!

OpenStudy (kamille):

2sin2x−3=0 doesnt have sollutions, right?

hartnn (hartnn):

correct.

hartnn (hartnn):

just solve cos 2x=0

OpenStudy (kamille):

thanks, hartnn!:)

hartnn (hartnn):

welcome :)